OpenStudy (anonymous):

Simplify. 17 – 6 × 10 ÷ 2 + 12 givin medals:)

OpenStudy (anonymous):

17 - 6 * 10 / 2 + 12 = 17 - (6 * 10 / 2) + 12 = 17 - 30 + 12 = 28 - 30 = -1 Therefore: 17 - 6 * 10 / 2 + 12 = -1

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

agradecimiento @nannaberry

OpenStudy (anonymous):

am i correct using PEMDAS P.arenthesis E.stimate M.mulitply D.ivide A.ddition S.ubtraction

OpenStudy (anonymous):

gratitude right?

OpenStudy (anonymous):

got it thanks @Starr_DynastyT

OpenStudy (solomonzelman):

lets do the operations in order. Multiplication and division THEN addition and subtraction \[17-\color{blue} { 6 \times 10 } \div 2+12\]\[17-\color{blue} { 60 } \div 2 +12\]NEXT\[17-\color{red} { 60 \div 2 } +12\]\[17-\color{red} { 30 } +12\]and FINALLY\[\color{green} { 17- 30 }+12\]\[\color{green} { - 13 }+12\]so,\[\color{goldenrod} { - 13+12=-1 }\]

OpenStudy (anonymous):

kbye i got it

OpenStudy (anonymous):

solomon did it right

OpenStudy (anonymous):

@nannaberry did it thats why i gave the medal df

OpenStudy (anonymous):

nooooooooooooo gd @nannaberry juss gave meh the answer

OpenStudy (anonymous):

gracias

OpenStudy (solomonzelman):

thank you in spanish?

OpenStudy (anonymous):

si

OpenStudy (anonymous):

sea lo que sea @nannaberry

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

PIE!!!

OpenStudy (anonymous):

im mucho lo de ese error, supongo que estaba buscando de forma diferente

OpenStudy (anonymous):

que está bien

OpenStudy (anonymous):

usted acaba de decir que está bien @nannaberry

OpenStudy (anonymous):

huh? i said it is ok

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

BLLLAAAAHHHH!!!

OpenStudy (anonymous):

RRRAAAAWWWRRRRR!!!!

OpenStudy (anonymous):

VVVVRRRRROOOOOOOOMMMM!!!!! BOOOOOOOMMMMM!!!!

OpenStudy (anonymous):

pow